# What One Cannot, Two Can: Two-Layer Transformers Provably Represent Induction Heads on Any-Order Markov Chains

### v1

@@ B.1 Three Layer MLP followed by the Second Attention Layer
$\textbf{W}_{\texttt{mlp}}^{(2)}$ $=\begin{bmatrix}\textbf{0}_{3\times 3}&\textbf{0}_{3\times S}&\textbf{0}_{3\times S}&\textbf{0}_{3\times S}&\textbf{0}_{3\times S}&\textbf{0}_{3\times S}&\textbf{0}_{3\times S}\\ \textbf{0}_{S\times 3}&{\textbf{0}_{S\times S}}&\textbf{0}_{S\times S}&\textbf{0}_{S\times S}&\textbf{0}_{S\times S}&\textbf{0}_{S\times S}&\textbf{0}_{S\times S}\\ \textbf{0}_{S\times 3}&\textbf{0}_{S\times S}&\textbf{0}_{S\times S}&\textbf{0}_{S\times S}&\textbf{0}_{S\times S}&\textbf{0}_{S\times S}&\textbf{0}_{S\times S}\\ \textbf{0}_{S\times 3}&{\frac{1}{C_{\texttt{attn}}}\cdot I_{S\times S}}&{-\frac{3^{k}}{C_{\texttt{attn}}}\cdot\textbf{A}^{T}_{S\times S}}&\textbf{0}_{S\times S}&{\frac{3}{C_{\texttt{attn}}}I_{S\times S}}&\textbf{0}_{S\times S}&\textbf{0}_{S\times S}\\ \textbf{0}_{S\times 3}&\textbf{0}_{S\times S}&\textbf{0}_{S\times S}&\textbf{0}_{S\times S}&\textbf{0}_{S\times S}&\textbf{0}_{S\times S}&\textbf{0}_{S\times S}\\ \textbf{0}_{S\times 3}&\textbf{0}_{S\times S}&\textbf{0}_{S\times S}&\textbf{0}_{S\times S}&{\textbf{0}_{S\times S}}&\textbf{0}_{S\times S}&\textbf{0}_{S\times S}\\ \textbf{0}_{S\times 3}&\textbf{0}_{S\times S}&\textbf{0}_{S\times S}&\textbf{0}_{S\times S}&\textbf{0}_{S\times S}&\textbf{0}_{S\times S}&\textbf{0}_{S\times S}\\ \end{bmatrix}$
@@
$\textbf{u}_{n}$ $=\frac{1}{C_{\texttt{attn}}}e_{x_{n}^{S}}+\frac{3}{C_{\texttt{attn}}}\textbf{v}_{n}-\frac{3^{k}}{C_{\texttt{attn}}}e^{S}_{x_{n-k}}$
@@ B.2 Two Layer MLP followed by the Second Attention Layer
$\textbf{W}_{\texttt{mlp}}^{(2)}$ $=\begin{bmatrix}{{I}_{3\times 3}}&\textbf{0}_{3\times S}&\textbf{0}_{3\times S}&\textbf{0}_{3\times S}&\textbf{0}_{3\times S}&\textbf{0}_{3\times S}&\textbf{0}_{3\times S}\\ \textbf{0}_{S\times 3}&{I_{S\times S}}&\textbf{0}_{S\times S}&\textbf{0}_{S\times S}&\textbf{0}_{S\times S}&\textbf{0}_{S\times S}&\textbf{0}_{S\times S}\\ \textbf{0}_{S\times 3}&{\frac{1}{C_{\texttt{attn}}}\cdot I_{S\times S}}&{-\frac{3^{k}}{C_{\texttt{attn}}}\cdot\textbf{A}^{T}_{S\times S}}&\textbf{0}_{S\times S}&{\frac{3}{C_{\texttt{attn}}}I_{S\times S}}&\textbf{0}_{S\times S}&\textbf{0}_{S\times S}\\ \textbf{0}_{S\times 3}&\textbf{0}_{S\times S}&\textbf{0}_{S\times S}&\textbf{0}_{S\times S}&\textbf{0}_{S\times S}&\textbf{0}_{S\times S}&\textbf{0}_{S\times S}\\ \textbf{0}_{S\times 3}&\textbf{0}_{S\times S}&\textbf{0}_{S\times S}&\textbf{0}_{S\times S}&{I_{S\times S}}&\textbf{0}_{S\times S}&\textbf{0}_{S\times S}\\ \textbf{0}_{S\times 3}&\textbf{0}_{S\times S}&\textbf{0}_{S\times S}&\textbf{0}_{S\times S}&\textbf{0}_{S\times S}&\textbf{0}_{S\times S}&\textbf{0}_{S\times S}\\ \textbf{0}_{S\times 3}&\textbf{0}_{S\times S}&\textbf{0}_{S\times S}&\textbf{0}_{S\times S}&\textbf{0}_{S\times S}&\textbf{0}_{S\times S}&\textbf{0}_{S\times S}\\ \end{bmatrix}$
@@
$\textbf{u}_{n}$ $=\frac{1}{C_{\texttt{attn}}}e_{x_{n}^{S}}+\frac{3}{C_{\texttt{attn}}}\textbf{v}_{n}-\frac{3^{k}}{C_{\texttt{attn}}}e^{S}_{x_{n-k}}$
@@ Appendix C Two-Layer Single-Head Gradient Analysis
As discussed in Sec. 5, we analyze a simplified model that retains only the key components of the two-layer architecture described in Thm. 4—specifically, the positional encodings and the attention layer—while treating the remaining components as approximately optimal. This reduction enables a tractable yet faithful reproduction of the in-context learning (ICL) behavior exhibited by the full model. Our approach is closely related to the reparameterization techniques used by Makkuva et al. [2024b] and Nichani et al. [2024] in the context of gradient flow and gradient descent analysis. Alternatively, it can be viewed as a form of optimal parameter initialization, in the spirit of Edelman et al. [2024] and Nichani et al. [2024]. We reiterate the assumptions and provide additional details whenever necessary.
@@
Following Nichani et al. [2024], we assume a prior distribution $\mathbb{P}_{\pi}(\cdot)$ over irreducible and aperiodic Markov transition kernels $\pi$, such that:
@@
We minimize the cross-entropy loss for next-token prediction Nichani et al. [2024]:
@@
Training algorithm. We denote the set of trainable parameters as $\theta=(\textbf{p},a_{2})$, where $\textbf{p}=[p_{0},p_{1},\ldots,p_{T}]^{T}$ represents the positional scalars and $a_{2}$ is the attention scalar. Following the approach of Nichani et al. [2024], we employ a two-stage training procedure. (i) In the first stage, we optimize the positional scalars p using gradient descent for $T_{1}$ steps with a learning rate $\eta_{1}$; (ii) in the second stage, we freeze p and train only $a_{2}$ for an additional $T_{2}$ steps using a distinct learning rate $\eta_{2}$. A key difference in our setup lies in the treatment of layer normalization: during the first stage, all layer norms in the MLP are omitted, whereas in the second stage, they are reintroduced. This stage-wise strategy facilitates a more tractable theoretical analysis. The training algorithm is:
@@
This section is organized as follows. We begin by presenting the simplified construction in App. C.1. Next, we demonstrate that during the first stage of training C.2, the model converges to the optimal attention map for the first layer—specifically, it learns to attend to the previous token, which is optimal under a first-order Markov chain. In the second stage of training C.3, the model sharpens its attention, ultimately converging asymptotically to a near-deterministic distribution over the desired token. Finally, we provide arguments for the result in the context of inductive generalization, utilizing results from Nichani et al. [2024]. We would like to note that our analysis for first-order processes is similar to that of Nichani et al. [2024], but the architecture is fundamentally different.
@@ C.2 Stage 1 Analysis
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
Hence, in the second stage of training, the inclusion of layer normalization does not significantly affect the outcome, provided the Markov chain is sufficiently long and the training duration is adequate, and hence, we will assume that $\textbf{v}_{i}/\|\textbf{v}_{i}\|\approx e_{x_{i-1}}$, which is the optimal attention map. We note that the gradient analysis for the first stage of training closely mirrors that of the first stage in Nichani et al. [2024], and is directly inspired by their approach.
@@ C.3 Stage 2 Analysis
We observe that, at this stage—assuming the attention map learned in the first stage is optimal by having the appropriate sequence length and training time—our simplified model aligns with the one presented in Nichani et al. [2024]. This alignment arises because, after the first stage of training, both models converge to the same optimal attention map in the first layer and subsequently train a single scalar parameter in the second attention layer. Therefore, we can directly invoke the theorems from Nichani et al. [2024] to analyze the convergence behavior of our model during the second stage of training. Accordingly, we state the relevant theorems below without proofs. Using the same set of assumptions outlined in Appendix C, we restate the key lemmas and results from Nichani et al. [2024] as follows.
Lemma 1 (Lemma D.8 in Nichani et al. [2024]).
@@
[…]
This clearly indicates that the negative gradient remains positive, causing $a_{2}$ to increase indefinitely under gradient descent. As a result, the attention mechanism progressively sharpens, effectively converging to a hard attention regime. Over time, this leads to convergence to the optimal $1$-gram estimator as the number of training iterations grows. For a thorough discussion of the time taken to converge, we refer readers to Nichani et al. [2024] (see Lemma D.10). Furthermore, given the equivalence between the models, once training is complete—assuming a sufficiently long sequence and adequate training duration—they are essentially identical. Therefore, we defer the proof of convergence of our model’s loss to the optimal loss (Theorem 5) to the corresponding argument in Nichani et al. [2024] (see D.7. Proof of Theorem 4.4). Finally, due to this model equivalence, and after the completion of both training stages, we again refer to Nichani et al. [2024] (see H.4. Proof of Theorem 4.5) for the proof of Theorem C.
@@ D.2 Convergence Analysis
Clearly, the above expression closely mirrors the reduced model presented in Nichani et al. [2024]. Since we have picked the softmax parameter $\kappa$ to be very small, we can apply a first-order Taylor approximation to simplify the expression. Let $\mathbf{1}_{n}=\left[1,1,\cdots,1\right]\in\mathbb{R}^{n+1}$. We also assume that $x_{n}=s_{1}$, $x_{n-1}=s_{2}$, $\ldots$, $x_{n-k}=s_{k}$.
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]

### v2

@@ B.1 Three Layer MLP followed by the Second Attention Layer
$\textbf{W}_{\texttt{mlp}}^{(2)}$ $=\begin{bmatrix}\textbf{0}_{3\times 3}&\textbf{0}_{3\times S}&\textbf{0}_{3\times S}&\textbf{0}_{3\times S}&\textbf{0}_{3\times S}&\textbf{0}_{3\times S}&\textbf{0}_{3\times S}\\ \textbf{0}_{S\times 3}&{\textbf{0}_{S\times S}}&\textbf{0}_{S\times S}&\textbf{0}_{S\times S}&\textbf{0}_{S\times S}&\textbf{0}_{S\times S}&\textbf{0}_{S\times S}\\ \textbf{0}_{S\times 3}&\textbf{0}_{S\times S}&\textbf{0}_{S\times S}&\textbf{0}_{S\times S}&\textbf{0}_{S\times S}&\textbf{0}_{S\times S}&\textbf{0}_{S\times S}\\ \textbf{0}_{S\times 3}&{\frac{1}{C_{\texttt{attn}}}\cdot I_{S\times S}}&{-\frac{3^{k}}{C_{\texttt{attn}}}\cdot\textbf{A}^{T}_{S\times S}}&\textbf{0}_{S\times S}&3\cdot{I_{S\times S}}&\textbf{0}_{S\times S}&\textbf{0}_{S\times S}\\ \textbf{0}_{S\times 3}&\textbf{0}_{S\times S}&\textbf{0}_{S\times S}&\textbf{0}_{S\times S}&\textbf{0}_{S\times S}&\textbf{0}_{S\times S}&\textbf{0}_{S\times S}\\ \textbf{0}_{S\times 3}&\textbf{0}_{S\times S}&\textbf{0}_{S\times S}&\textbf{0}_{S\times S}&{\textbf{0}_{S\times S}}&\textbf{0}_{S\times S}&\textbf{0}_{S\times S}\\ \textbf{0}_{S\times 3}&\textbf{0}_{S\times S}&\textbf{0}_{S\times S}&\textbf{0}_{S\times S}&\textbf{0}_{S\times S}&\textbf{0}_{S\times S}&\textbf{0}_{S\times S}\\ \end{bmatrix}$
@@
$\textbf{u}_{n}$ $=\frac{1}{C_{\texttt{attn}}}e_{x_{n}^{S}}+3\textbf{v}_{n}-\frac{3^{k}}{C_{\texttt{attn}}}e^{S}_{x_{n-k}}$
@@ B.2 Two Layer MLP followed by the Second Attention Layer
$\textbf{W}_{\texttt{mlp}}^{(2)}$ $=\begin{bmatrix}{{I}_{3\times 3}}&\textbf{0}_{3\times S}&\textbf{0}_{3\times S}&\textbf{0}_{3\times S}&\textbf{0}_{3\times S}&\textbf{0}_{3\times S}&\textbf{0}_{3\times S}\\ \textbf{0}_{S\times 3}&{I_{S\times S}}&\textbf{0}_{S\times S}&\textbf{0}_{S\times S}&\textbf{0}_{S\times S}&\textbf{0}_{S\times S}&\textbf{0}_{S\times S}\\ \textbf{0}_{S\times 3}&{\frac{1}{C_{\texttt{attn}}}\cdot I_{S\times S}}&{-\frac{3^{k}}{C_{\texttt{attn}}}\cdot\textbf{A}^{T}_{S\times S}}&\textbf{0}_{S\times S}&3\cdot{I_{S\times S}}&\textbf{0}_{S\times S}&\textbf{0}_{S\times S}\\ \textbf{0}_{S\times 3}&\textbf{0}_{S\times S}&\textbf{0}_{S\times S}&\textbf{0}_{S\times S}&\textbf{0}_{S\times S}&\textbf{0}_{S\times S}&\textbf{0}_{S\times S}\\ \textbf{0}_{S\times 3}&\textbf{0}_{S\times S}&\textbf{0}_{S\times S}&\textbf{0}_{S\times S}&{I_{S\times S}}&\textbf{0}_{S\times S}&\textbf{0}_{S\times S}\\ \textbf{0}_{S\times 3}&\textbf{0}_{S\times S}&\textbf{0}_{S\times S}&\textbf{0}_{S\times S}&\textbf{0}_{S\times S}&\textbf{0}_{S\times S}&\textbf{0}_{S\times S}\\ \textbf{0}_{S\times 3}&\textbf{0}_{S\times S}&\textbf{0}_{S\times S}&\textbf{0}_{S\times S}&\textbf{0}_{S\times S}&\textbf{0}_{S\times S}&\textbf{0}_{S\times S}\\ \end{bmatrix}$
@@
$\textbf{u}_{n}$ $=\frac{1}{C_{\texttt{attn}}}e_{x_{n}^{S}}+3\textbf{v}_{n}-\frac{3^{k}}{C_{\texttt{attn}}}e^{S}_{x_{n-k}}$
@@ Appendix C Two-Layer Single-Head Gradient Analysis
As discussed in Sec. 5, we analyze a simplified model that retains only the key components of the two-layer architecture described in Thm. 4—specifically, the positional encodings and the attention layer—while treating the remaining components as approximately optimal. This reduction enables a tractable yet faithful reproduction of the in-context learning (ICL) behavior exhibited by the full model. Our approach is closely related to the reparameterization techniques used by makkuva2024local and nichani2024how in the context of gradient flow and gradient descent analysis. Alternatively, it can be viewed as a form of optimal parameter initialization, in the spirit of edelman2024evolution and nichani2024how. We reiterate the assumptions and provide additional details whenever necessary.
@@
Following nichani2024how, we assume a prior distribution $\mathbb{P}_{\pi}(\cdot)$ over irreducible and aperiodic Markov transition kernels $\pi$, such that:
@@
We minimize the cross-entropy loss for next-token prediction [nichani2024how]:
@@
Training algorithm. We denote the set of trainable parameters as $\theta=(\textbf{p},a_{2})$, where $\textbf{p}=[p_{0},p_{1},\ldots,p_{T}]^{T}$ represents the positional scalars and $a_{2}$ is the attention scalar. Following the approach of [nichani2024how], we employ a two-stage training procedure. (i) In the first stage, we optimize the positional scalars p using gradient descent for $T_{1}$ steps with a learning rate $\eta_{1}$; (ii) in the second stage, we freeze p and train only $a_{2}$ for an additional $T_{2}$ steps using a distinct learning rate $\eta_{2}$. A key difference in our setup lies in the treatment of layer normalization: during the first stage, all layer norms in the MLP are omitted, whereas in the second stage, they are reintroduced. This stage-wise strategy facilitates a more tractable theoretical analysis. The training algorithm is:
@@
This section is organized as follows. We begin by presenting the simplified construction in App. C.1. Next, we demonstrate that during the first stage of training C.2, the model converges to the optimal attention map for the first layer—specifically, it learns to attend to the previous token, which is optimal under a first-order Markov chain. In the second stage of training C.3, the model sharpens its attention, ultimately converging asymptotically to a near-deterministic distribution over the desired token. Finally, we provide arguments for the result in the context of inductive generalization, utilizing results from nichani2024how. We would like to note that our analysis for first-order processes is similar to that of nichani2024how, but the architecture is fundamentally different.
@@ C.2 Stage 1 Analysis
$\frac{\mathrm{d}L(\theta)}{\mathrm{d}p_{m}}$ $=-\frac{1}{S}\mathbb{E}_{\pi\sim\mathbb{P}(\pi),X_{n}}\left[{\sum_{s^{\prime},s}\pi\left({s_{n+1}=s^{\prime}|s}\right)\frac{\mathrm{d}\log\left({\operatorname{logit}_{s^{\prime}_{s},n}}\right)}{\mathrm{d}p_{m}}}\right]$
@@
$\frac{\mathrm{d}\log\left(\operatorname{logit}_{s^{\prime}_{s},n}\right)}{\mathrm{d}p_{m}}$ $=\frac{a_{2,0}}{S\cdot n\cdot\mu_{\pi}(s^{\prime})}\cdot\sum_{i=m}^{n}\frac{\exp(p_{m})}{\sum_{j=0}^{i}\exp(p_{j})}\cdot\left[q_{i,m}-\sum_{j=0}^{i}\frac{\exp(p_{j}\cdot a_{1})}{\sum_{j=0}^{i}\exp(p_{j}\cdot a_{1})}q_{i,j}\right]$
@@
$\frac{\mathrm{d}L(\theta)}{\mathrm{d}p_{m}}$ $=-\frac{a_{2,0}}{S\cdot kn}\cdot\mathbb{E}_{\pi\sim\mathbb{P}(\pi)}\Bigg[\sum_{i\geq m}^{n}\sum_{s^{\prime},s}\frac{\pi(s^{\prime}|s)}{\mu_{\pi}(s^{\prime})}\cdot\frac{\exp(p_{m})}{\sum_{j=0}^{i}\exp(p_{j})}$
@@
$\frac{\mathrm{d}L(\theta)}{\mathrm{d}p_{m}}$ $=-\frac{a_{2,0}}{S\cdot kn}\cdot\mathbb{E}_{\pi\sim\mathbb{P}(\pi)}\Bigg[\sum_{i\geq m}^{n}\frac{\exp(p_{m})}{\sum_{j=0}^{i}\exp(p_{j})}g_{i,m}(\pi)-\sum_{i\geq m}^{n}\sum_{j=0}^{i}\frac{\exp(p_{j})\,\exp(p_{m})}{\left(\sum_{j=0}^{i}\exp(p_{j})\right)^{2}}g_{i,j}(\pi)\Bigg]$
@@
$\frac{\mathrm{d}L(\theta)}{\mathrm{d}p_{m}}\bigg|_{i}$ $=-\frac{a_{2,0}}{S\cdot kn}\cdot\mathbb{E}_{\pi\sim\mathbb{P}(\pi)}\left[{\left({\frac{\exp(p_{m})}{\sum_{j=0}^{i}\exp(p_{j})}g_{i,m}(\pi)-\sum_{j=0}^{i}\frac{\exp(p_{j})\,\exp(p_{m})}{\left(\sum_{j=0}^{i}\exp(p_{j})\right)^{2}}g_{i,j}(\pi)}\right)}\right]$
@@
$\left.\frac{\mathrm{d}L(\theta)}{\mathrm{d}p_{1}}\right|_{i}$ $>\left.\frac{\mathrm{d}L(\theta)}{\mathrm{d}p_{k}}\right|_{i}\quad\forall\,k\in\{0,2,\ldots,n\}$
$\left.\frac{\mathrm{d}L(\theta)}{\mathrm{d}p_{1}}\right|_{i}$ $>0$
In this section, we will show that given the prior assumptions, this is indeed true. Towards this, we denote $d_{g_{m,i}}=\left.\frac{\mathrm{d}L(\theta)}{\mathrm{d}p_{m}}\right|_{i}$. With this notation, we can express:
@@
Hence, in the second stage of training, the inclusion of layer normalization does not significantly affect the outcome, provided the Markov chain is sufficiently long and the training duration is adequate, and hence, we will assume that $\textbf{v}_{i}/\|\textbf{v}_{i}\|\approx e_{x_{i-1}}$, which is the optimal attention map. We note that the gradient analysis for the first stage of training closely mirrors that of the first stage in nichani2024how, and is directly inspired by their approach.
@@ C.3 Stage 2 Analysis
We observe that, at this stage—assuming the attention map learned in the first stage is optimal by having the appropriate sequence length and training time—our simplified model aligns with the one presented in nichani2024how. This alignment arises because, after the first stage of training, both models converge to the same optimal attention map in the first layer and subsequently train a single scalar parameter in the second attention layer. Therefore, we can directly invoke the theorems from nichani2024how to analyze the convergence behavior of our model during the second stage of training. Accordingly, we state the relevant theorems below without proofs. Using the same set of assumptions outlined in Appendix C, we restate the key lemmas and results from nichani2024how as follows.
Lemma 1 (Lemma D.8 in nichani2024how).
@@
$1\geq-\frac{\mathrm{d}L(\theta)}{\mathrm{d}a_{2}}\geq\frac{1}{4}\gamma^{8}S^{-6}e^{-2a_{2}}>0.$
This clearly indicates that the negative gradient remains positive, causing $a_{2}$ to increase indefinitely under gradient descent. As a result, the attention mechanism progressively sharpens, effectively converging to a hard attention regime. Over time, this leads to convergence to the optimal $1$-gram estimator as the number of training iterations grows. For a thorough discussion of the time taken to converge, we refer readers to nichani2024how (see Lemma D.10). Furthermore, given the equivalence between the models, once training is complete—assuming a sufficiently long sequence and adequate training duration—they are essentially identical. Therefore, we defer the proof of convergence of our model’s loss to the optimal loss (Theorem 5) to the corresponding argument in nichani2024how (see D.7. Proof of Theorem 4.4). Finally, due to this model equivalence, and after the completion of both training stages, we again refer to nichani2024how (see H.4. Proof of Theorem 4.5) for the proof of Theorem C.
@@ D.2 Convergence Analysis
Clearly, the above expression closely mirrors the reduced model presented in nichani2024how. Since we have picked the softmax parameter $\kappa$ to be very small, we can apply a first-order Taylor approximation to simplify the expression. Let $\mathbf{1}_{n}=\left[1,1,\cdots,1\right]\in\mathbb{R}^{n+1}$. We also assume that $x_{n}=s_{1}$, $x_{n-1}=s_{2}$, $\ldots$, $x_{n-k}=s_{k}$.
@@
$\frac{\mathrm{d}L(\theta)}{\mathrm{d}p_{m}}$ $=-\frac{1}{|\mathcal{S}|^{k}}\mathbb{E}_{\pi\sim\mathbb{P}(\pi),\,s_{n}\sim\pi}\left[\sum_{s^{\prime},s_{1},\ldots,s_{k}}\pi(s_{n+1}=s^{\prime}\mid s_{1},\ldots,s_{k})\cdot\frac{\mathrm{d}\log\left(\operatorname{logit}_{s^{\prime}_{s},n}\right)}{\mathrm{d}p_{m}}\right]$
@@
$\frac{\mathrm{d}\log\left(\operatorname{logit}_{s^{\prime}_{s},n}\right)}{\mathrm{d}p_{m}}$ $=\frac{\kappa}{|\mathcal{S}|^{k}\cdot kn\cdot\mu_{\pi}(s^{\prime})}\cdot\sum_{l=1}^{k}\sum_{i=m}^{n}\frac{\exp(p_{m})}{\sum_{j=0}^{i}\exp(p_{j})}\cdot\left[q_{i,m,l}-\sum_{j=0}^{i}\frac{\exp(p_{j})}{\sum_{j=0}^{i}\exp(p_{j})}q_{i,j,l}\right]$
@@
$\frac{\mathrm{d}L(\theta)}{\mathrm{d}p_{m}}$ $=-\frac{\kappa}{|\mathcal{S}|^{k}\cdot kn}\cdot\mathbb{E}_{\pi\sim\mathbb{P}(\pi)}\Bigg[\sum_{i\geq m}^{n}\sum_{l=1}^{k}\sum_{s^{\prime},s_{1},\ldots,s_{k}}\frac{\pi(s^{\prime}|s_{1},\ldots,s_{k})}{\mu_{\pi}(s^{\prime})}\cdot\frac{\exp(p_{m})}{\sum_{j=0}^{i}\exp(p_{j})}$
@@
$\frac{\mathrm{d}L(\theta)}{\mathrm{d}p_{m}}\bigg|_{i}$ $=-\frac{\kappa}{|\mathcal{S}|^{k}\cdot kn}\cdot\mathbb{E}_{\pi\sim\mathbb{P}(\pi)}\Bigg[\sum_{l=1}^{k}\sum_{s^{\prime},s_{1},\ldots,s_{k}}\frac{\pi(s^{\prime}|s_{1},\ldots,s_{k})}{\mu_{\pi}(s^{\prime})}\cdot\frac{\exp(p_{m})}{\sum_{j=0}^{i}\exp(p_{j})}$
@@
$\frac{\mathrm{d}L(\theta)}{\mathrm{d}p_{m}}\bigg|_{i}$ $=-\frac{\kappa}{|\mathcal{S}|^{k}\cdot kn}\cdot\mathbb{E}_{\pi\sim\mathbb{P}(\pi)}\left[{\left({\frac{\exp(p_{m})}{\sum_{j=0}^{i}\exp(p_{j})}g_{i,m,}-\sum_{j=0}^{i}\frac{\exp(p_{j})\,\exp(p_{m})}{\left(\sum_{j=0}^{i}\exp(p_{j})\right)^{2}}g_{i,j}}\right)}\right]$